(TCO 5) Explain the ventilation cycle and how it relates to the respiratory rate. If a patient breathes at a rate of 14 breaths per minute and has an inspiratory time of 2.5 seconds, what is the expiratory time?
Answer #1

The change from inspiration to expiration in the mechanically ventilated breath is termed as cycling. Mechanical ventilation optimization and synchronization with patients breathing is very important to maintain efficient respiration. The mechanical ventilator breath can be divided into inspiratory phase and expiratory phase. At the end of inspiration, gas flow ceases, and the breath switches exhalation. This transition of the inspiratory phase to the expiratory phase is termed “cycling.” The cycling depends on four variables : pressure, time, volume and flow.

Total cycle time= 60/respiratory rate = 60/14 = 4.3 sec

Now the inspiratory time given is 2.5 sec

So the expiratiry time = total cycle time - inspiratory time

Expiratory time = 4.3 - 2.5 = 1.8sec